OpenStudy (anonymous):

what is the force of a truck that has a mass of 1000 kg moving at 80 km/h with an acceleration of 0 m/s(squared)?

OpenStudy (anonymous):

Well if we use the equation F=change in (mv)/change in time we get F= 1000*80000/1 so we get 8 x10^7 N. We have 80000 because we need the speed in m per hour than seeing as though the time is 1 hour we use change in time =1

OpenStudy (anonymous):

Force = Mass times acceleration (F=MA) If the truck isn't accelerating. Then it has no Force. It does have momentum, mass times velocity (MV).

OpenStudy (anonymous):

Yes sorry I forgot to mention, that would be the force applied to an object if it were to decelerate the truck to a stop. So at the moment it has momentum but no resultant force

OpenStudy (anonymous):

Momentum = Inertial Mass X Velocity P = mv P = (1000kg) x (22.2m/s) P = 22,200 kg-m/s

OpenStudy (anonymous):

F = ma. Since a = 0 m/(s^2) , therefore F = 0 N.

OpenStudy (anonymous):

SINCE THE CAR IS MOVING WITH A CONSTATNT VELOCITY THE SUMMITION OF EXTERNAL FORCES ON THE CAR IS ZERO.BUT THE MOTOR OF THE CAR IS APPLYING A FORCE EQUAL TO THE FORCE OF THE RESISTANCE OF AIR ,ROAL ,... BUT SUMMITION OF EXTERNAL FORCES IS ZERO

OpenStudy (anonymous):

The truck has 0 horizontal force given that it is not accelerating. f=ma --> f= 1000 x 0 = 0 HOWEVER, the truck is exerting a force of 9810 N down, due to the gravity.

OpenStudy (anonymous):

you guys need to be careful. The "net force" on the truck is zero.